OpenStudy (anonymous):

A relation R is defined on the real numbers by aRb if a-b is an integer. Prove that R is an equivalence relation and determine the equivalence class of 1/2

OpenStudy (anonymous):

You could try this type of argument, but I'm questioning my own correctness... if a and b and c must be real, then if d, e, q, r and n are integers: \[\frac{d}{e} - \frac{q}{r} = n\] where a = c/d and b = q/r, which implies: dr - qe = ner .... which are also all integers. Then the following is true for all integer pairs Int:Int, if a = d:r and b = q:e and ( c = d:r or q:e not equal with a or b respectively ) (a, a), (b, b), (c, c) are all members of R (reflexive) (a, b), (b,a) [where a or b are interchangable with c] are both members of R (symmetric) (a, b) , (b, c) and (a,c) are all members of R (transitive)

OpenStudy (anonymous):

anyone have any other ideas?